# Machine-Learning-Enhanced Optimization of Noise-Resilient Variational Quantum Eigensolvers

### v1

@@ 3.1 VQE Kernel
A GP needs to be equipped with a kernel (covariance) function. While there are many types of such functions, one popular example is the Gaussian-RBF kernel [15], which encodes into the GP the belief that input vectors that are close in the input space strongly correlate compared to points that are far apart, with the correlation strength decaying exponentially with the distance. Although this might be a reasonable assumption, an exponential drop in correlation, like in the Gaussian-RBF kernel, may not be representative of the underlying features of every problem. For this reason, wisely choosing the kernel for a GP regression task can substantially affect the performance. In Ref. [9] the authors proposed the so-called VQE kernel:
@@
It was shown that the VQE kernel can be decomposed as
$k^{\mathrm{VQE}}\left(\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\right)=\boldsymbol{\phi}(\boldsymbol{\theta})^{\top}\boldsymbol{\phi}(\boldsymbol{\theta}^{\prime}),\quad\textrm{with}\quad\boldsymbol{\phi}(\boldsymbol{\theta})=\sigma_{0}\left(1+\gamma^{2}\right)^{-D/2}\left[\otimes_{d=1}^{D}\left(\gamma,\cos\theta_{d},\sin\theta_{d}\right)^{\top}\right]$ (11)
where the feature maps $\boldsymbol{\phi}(\boldsymbol{\theta})$ fulfill the same VQE functional form derived in Eq. 4. Since a GP regression model is only defined in terms of inner products of the feature maps, one can replace all of these occurrences with the kernel described above – a procedure often known as kernel trick [15].
Using the VQE kernel function instead of, e.g., the RBF-kernel, ensures that the GP samples functions that match the required functional form from Eq. 4. Consequently, as described in Eq. 6, when optimizing on a one-dimensional subspace, i.e., changing only one parameter at a time, three (noiseless) observed points allow to uniquely identify the target function [7] over the entire one-dimensional subspace. However, in the presence of statistical uncertainty due to repeated measurements (shots), the target function in the subspace can be identified up to some statistical uncertainty. Therefore, the GP will be able to provide energy estimates, and their corresponding uncertainties, for new measurements in the subspace.
Furthermore, it has been shown that the decomposition of the VQE kernel in Eq. 11 can be generalized so that each entry of $\boldsymbol{\theta}$ may parametrize multiple gates at once [9]. This property becomes crucial if the Hamiltonian is invariant under certain symmetries, which results in multiple parameters being identical and requiring simultaneous rotation.
3.2 EMICoRe: Expected Maximum Improvement over Confident Regions
To perform Bayesian Optimization (BO) in combination with GPR, one should pick a suitable acquisition function to identify the most promising points to measure at each optimization step. A novel acquisition function termed Expected Maximum Improvement over Confident Regions (EMICoRe) was introduced in the context of GPR for VQEs [9]. This acquisition function relies on the idea of Confident Region (CoRe) mathematically defined as
@@ 3.2 EMICoRe: Expected Maximum Improvement over Confident Regions
The CoRe describes all points for which the predictive covariance $s_{\boldsymbol{\Theta}}(\cdot,\cdot)$ is smaller or equal to a threshold $\kappa^{2}$, which is a hyperparameter of the model. For sufficiently small $\kappa$, an appropriate kernel function, and a weak prior, i.e., enough previously observed points, every point in the CoRe can be regarded as “already observed” because the uncertainty of the GP is small enough. Leveraging the concept of CoRe, Nicoli et al. [9] proposed the following acquisition function
$a^{\mathrm{EMICoRe}}\left(\boldsymbol{\Theta}^{\prime}\right)=\frac{1}{M}\left\langle\max\left(0,\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\boldsymbol{\Theta}}}f(\boldsymbol{\theta})-\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\widetilde{\boldsymbol{\Theta}}}}f(\boldsymbol{\theta})\right)\right\rangle_{p(f(\cdot)\mid\boldsymbol{\Theta},\boldsymbol{y})},$ (13)
where $\widetilde{\boldsymbol{\Theta}}=\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\in\mathcal{X}^{N+M}$ denotes the augmented training set with the new input points $\boldsymbol{\Theta}^{\prime}\in\mathcal{X}^{M}$, and the previously observed ones $\boldsymbol{\Theta}^{\prime}\in\mathcal{X}^{N}$.
The function $\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\boldsymbol{\Theta}}}f(\boldsymbol{\theta})$ is the estimated minimum of the CoRe, identified by all previously observed points $\boldsymbol{\Theta}$. Such value is drawn from the GP’s posterior distribution and identifies the estimated minimum energy for any choice of circuit parameters previously observed. The other term in Eq. 13, i.e., $\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\widetilde{\boldsymbol{\Theta}}}}f(\boldsymbol{\theta})$, is the estimated minimum over the augmented training set $\widetilde{\boldsymbol{\Theta}}$. The new input points $\boldsymbol{\Theta}^{\prime}\in\mathcal{X}^{M}$ are treated as if they were observed. In fact, when a point is in the CoRe, the GP’s uncertainty is smaller than a given threshold hence allowing to make the reasonable assumption that the GP is highly confident in a close range from that point. Therefore, one could implicitly assume this point as known, i.e., as if it has been observed (even though no measurement on the quantum computer has been performed).
3.2.1 EMICoRe
Since the NFT algorithm is not designed to deal with the statistical noise arising from measuring on a quantum computer, Nicoli et al. [9] proposed to combine NFT with GP regression and BO, thus merging all the tools previously introduced in Sections 2.2, 3.1 and 3.2.
The proposed algorithm builds on the structure of the NFT algorithm but improves it with two crucial modifications:
@@ 3.2.1 EMICoRe
    Learnable shifts: Instead of taking a fixed shift $\alpha$ to identify two new (equidistant) candidate points, one uses the EMICoRe acquisition function from Section 3.2, to identify the most informative points to measure next.
@@
    Re-use of previous measurements: For every iteration the estimation for the minimum on the one-dimensional subspace (using Least Square Regression) is performed using the posterior mean of the GP. This implies that the GP will give better estimates as more measurements are made on the quantum computer and used to train the surrogate model.
@@
These modifications make the proposed algorithm very different from NFT, where the estimation of the minimum on the subspace solely relies on the previous best point and the fixed shifts for the new measurements. Both modifications aim at making EMICoRe a more flexible yet resilient algorithm with respect to the noise of the quantum computer. The EMICoRe algorithm starts with $T_{\textrm{NFT}}$ steps of the baseline NFT algorithm, necessary to collect sufficient measurements to instantiate an informative GP to make predictions. After this initial phase, the EMICoRe algorithm starts learning the shifts, for each step, by using the previous measurements and storing the new measurements for future predictions. For a detailed description of the EMICoRe algorithm and its subroutine, we refer to the original paper [9].
@@
Figure 2 visualizes the algorithmic procedure of the full EMICoRe algorithm. The objective function $E^{\ast}(\boldsymbol{\theta})$ (true energy) is represented by the red lines, while the red shaded area indicates the standard deviation of an measurements on the quantum computer. The blue lines indicate the predictive mean of the GP $\mu_{\boldsymbol{\Theta}}(\cdot)$, while the shaded area denotes its predictive uncertainty $\sqrt{s_{\boldsymbol{\Theta}}(\cdot,\cdot)}$, i.e., one standard deviation. The figure should be read from top left to bottom right following the black arrow.
@@
The top left plot shows the optimization in the direction $\theta_{d}$ for $d\in\{1,\ldots,D\}$ [^2] Note the direction to be optimized is chosen sequentially until $d=D$, after which the algorithm restarts from $d=0$.. The quantity ${\boldsymbol{\hat{\theta}}}^{(t)}$ denotes the set of parameters which give the lowest energy $E^{\ast}({\boldsymbol{\hat{\theta}}}^{(t)})$ after the $t-$th optimization step. Note that at this point the GP has access to a dataset of previous measurements $\{\boldsymbol{\Theta}^{(t)},\boldsymbol{\tilde{E}}^{(t)}\}$. This explains why the blue line, and the corresponding uncertainty, already give a good prediction of the red curve along direction $d$, even though no optimization has been performed yet.
@@
The EMICoRe subroutine is now responsible for finding the promising points to be measured next, i.e., $\boldsymbol{\Theta}^{\prime}$ in Eq. 7 with the shifts $\alpha$ being different for $\theta_{1}^{\prime}$ and $\theta_{2}^{\prime}$. First, the subroutine divides the subspace into a grid of equally spaced points, as represented by the grid lines at the bottom of the second plot. Second, it prepares all possible combinations of candidate pairs $\boldsymbol{\Theta}^{\prime}_{k}=\{\boldsymbol{\theta}^{\prime}_{k,2},\boldsymbol{\theta}^{\prime}_{k,2}\}$. These points are added to the base-GP by pairs so that a discretized CoRe for each pair can be computed (orange vertical lines). The subroutine then computes the acquisition function for each candidate pair $\boldsymbol{\Theta}^{\prime}_{k}$ using the respective CoRe. The pair, which maximizes the acquisition function, see Eq. 13, is returned as $\boldsymbol{\Theta}^{\prime}$. According to the algorithm, this pair represents the points which are the most promising to be observed next. The fifth plot shows the measurements on the quantum computer at the suggested points, which are indicated by the blue crosses. The GP is finally updated using the new measurements— note the change of the blue curve and the shaded area in the plot. Finally, the GP’s posterior mean is computed for three equidistant points on the 1D-subspace. These points are then used to fit a cosine function according to the functional form of the subspace in Eq. 6. This fitted cosine is then used to obtain the analytic minimum of the 1D-subspace.
@@
We recall that because of the special form of the VQE-kernel and the design of the EMICoRe algorithm, the functional sampled from the GP, i.e, the posterior mean, follows the same functional of Eq. 4. Therefore, fitting a sinusoidal function to the GP’s posterior is a valid assumption.
@@ 4 Numerical Experiments
In the following, we perform some numerical experiments to demonstrate the performance of EMICoRe in the presence of hardware noise. Our experiments focus on the quantum Heisenberg Hamiltonian, a typical Hamiltonian for benchmarking the performance of VQEs [16]
@@
where the index of the Pauli matrices $\{X_{j},Y_{j},Z_{j}\}$ refers to the qubit that the Pauli gate is applied to, and open boundary conditions are considered. In this work, we chose the couplings to be at the critical point with $(J_{X},J_{Y},J_{Z})=(-1,0,0)$ and $(h_{X},h_{Y},h_{Z})=(0,0,-1)$. The resulting Hamiltonian is also known as the Ising Hamiltonian at criticality.
In order to assess the performance of the VQE during the optimization, we mainly look at two metrics: energy and fidelity. The former is the lowest energy reached by the VQE during the optimization process, i.e., the expectation value of the Hamiltonian with respect to the variational state. The fidelity, measures the overlap between the variational (VQE) state and the true ground-state obtained with exact diagonalization. We refer to [9] for more details on the metrics we used. The ground-state energy $\underline{E}^{\ast}=\langle\psi_{\mathrm{GS}}|H|\psi_{\mathrm{GS}}\rangle$ together with the corresponding ground-state wave function $|\psi_{\mathrm{GS}}\rangle$ can be obtained in our case, because the size of the systems we tackle is sufficiently small. This is important as it gives us access to a ground truth for benchmarking different algorithms, i.e., NFT vs EMICoRe. To test the stability of the optimization with respect to initialization, we initialize several independent runs with different seeds. Specifically, we use $50$ runs for the experiments without Hardware noise and $10$ runs for the experiments with Hardware noise, see Section 4.1 for more details. In all plots we always report the median (50th percentile) as a solid line and the 25th and 75th percentiles being the lower and upper bounds respectively (shaded regions around the solid line). The results are plotted as a function of the total number of measurements performed on the quantum computer.
For all simulations, we use an $L$-layered and fully connected Efficient SU(2) circuit with $D=2(L+1)\cdot Q$ total angular parameters. The simulations of the quantum computer are performed with the Qiskit [17] library, while the Python implementation of the EMICoRe algorithm has been adapted from the original code available on GitHub [18].
All the parameters for EMICoRe are identical to the parameters chosen for the simulation of Fig. 7 in the original paper [9]. As noted in Footnote 1, we used $\pi/3$ as the shift for NFT. Moreover, in contrast to the original paper [7], we did not include a stabilization step, as our results showed that its use often result in poorer performance. We leave further investigation of this phenomenon to future work.
@@
Figure 3: Energy (left) and Fidelity (right) for NFT sequential (green) and EMICoRe (red). We show results for the critical Ising Hamiltonian using $Q=5$ qubits and $L=3$ layers, for an Efficient SU(2) quantum circuit, and $N_{\textrm{shots}}=1024$ measurement shots. The top row shows runs with simulated hardware noise and no error mitigation, while in the second and third rows results for classical simulations of noisy hardware are shown, with TREX and ZNE error mitigation scheme respectively. The statistics, e.g., the median (solid line), $25^{\mathrm{th}}$ and $75^{\mathrm{th}}$ percentiles (colored regions surrounding the median), are obtained with $50$ and $10$ independent seeded runs for experiments in Fig. 3(a) and Fig. 3(b), respectively. A trial density showing the distribution of the runs is shown right next to each plot.

### v2

@@ 3.1 VQE Kernel
A GP needs to be equipped with a kernel (covariance) function. While there are many types of such functions, one popular example is the Gaussian-RBF kernel [15]. This kernel encodes the assumption that input vectors that are close in the input space exhibit stronger correlations compared to those that are distant, with the correlation strength decaying exponentially as a function of the distance. While this may be a reasonable assumption, an exponential decay in correlation, as seen in the Gaussian-RBF kernel, may not accurately reflect the underlying features of every problem. For this reason, the choice of kernel in a GPR task can significantly affect performance. In Ref. [9] the authors proposed the so-called VQE kernel:
@@
In [9], it was shown that the VQE kernel can be decomposed as
$k^{\mathrm{VQE}}\left(\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\right)=\boldsymbol{\phi}(\boldsymbol{\theta})^{\top}\boldsymbol{\phi}(\boldsymbol{\theta}^{\prime}),\quad\textrm{with}\quad\boldsymbol{\phi}(\boldsymbol{\theta})=\sigma_{0}\left(1+\gamma^{2}\right)^{-D/2}\left[\otimes_{d=1}^{D}\left(\gamma,\cos\theta_{d},\sin\theta_{d}\right)^{\top}\right],$ (12)
where the feature maps $\boldsymbol{\phi}(\boldsymbol{\theta})$ fulfill the same VQE functional form as derived in Eq. 4. Since a GP regression model is only defined in terms of inner products of the feature maps, one just requires access to the kernel form without explicitly specifying $\boldsymbol{\phi}$ —a procedure known as kernel trick [15].
By using the VQE kernel instead of, for example, the Gaussian-RBF kernel, the GP is ensured to sample functions that align with the required functional form from Eq. 4. Consequently, as described in Eq. 6, when optimizing over a one-dimensional subspace—i.e., varying only one parameter at a time—three noiseless observed points are sufficient to uniquely identify the target function [7] across the entire subspace. However, in the presence of measurement shot noise, the target function in the subspace can only be identified up to some statistical uncertainty. Therefore, the GP provides not only energy estimates but also their corresponding uncertainties for new measurements in the subspace.
Furthermore, it has been shown that the decomposition of the VQE kernel in Eq. 12 can be generalized so that each entry of $\boldsymbol{\theta}$ may parametrize multiple gates at once [9]. This property becomes crucial if the Hamiltonian is invariant under specific symmetries, leading to multiple parameters being identical and requiring simultaneous updates.
3.2 Expected Maximum Improvement over Confident Regions
To perform BO in combination with GPR, one should pick a suitable acquisition function to identify the most promising points to measure at each optimization step. A novel acquisition function termed Expected Maximum Improvement over Confident Regions (EMICoRe) was introduced in the context of GPR for VQEs [9]. This acquisition function relies on the idea of Confident Region (CoRe) mathematically defined as
@@ 3.2 Expected Maximum Improvement over Confident Regions
The CoRe describes all points for which the predictive covariance $s_{\boldsymbol{\Theta}}(\cdot,\cdot)$ is smaller than or equal to a threshold $\kappa^{2}$, which is a hyperparameter of the model. For sufficiently small $\kappa$, an appropriate kernel function, and a weak prior—i.e., enough previously observed points—every point in the CoRe can be regarded as “already observed” due to the sufficiently small uncertainty of the GP. Leveraging the concept of CoRe, Nicoli et al. [9] proposed the following acquisition function:
$a^{\mathrm{EMICoRe}}\left(\boldsymbol{\Theta}^{\prime}\right)=\frac{1}{M}\left\langle\max\left(0,\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\boldsymbol{\Theta}}}E(\boldsymbol{\theta})-\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\widetilde{\boldsymbol{\Theta}}}}E(\boldsymbol{\theta})\right)\right\rangle_{p(E(\cdot)\mid\boldsymbol{\Theta},\boldsymbol{\tilde{E}})},$ (14)
where $\widetilde{\boldsymbol{\Theta}}=\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\in\mathcal{X}^{N+M}$ denotes the augmented training set with the new input points $\boldsymbol{\Theta}^{\prime}\in\mathcal{X}^{M}$ and the previously observed ones $\boldsymbol{\Theta}\in\mathcal{X}^{N}$.
[…]
3.2.1 The EMICoRe algorithm
To address the challenges posed by noisy measurements on quantum hardware, Nicoli et al. [9] proposed to combine NFT with GP regression and BO, thus merging all the tools previously introduced in Sections 2.2, 3.1 and 3.2.
The proposed EMICoRe algorithm builds upon the structure of the NFT algorithm, incorporating two key modifications:
@@ 3.2.1 The EMICoRe algorithm
    Learnable shifts: Instead of taking a fixed shift $\alpha$ to identify two new (equidistant) candidate points, see Eq. 7, the EMICoRe algorithm uses the EMICoRe acquisition function from Section 3.2 to identify the most informative points to measure next.
@@
    Re-use of previous measurements: At each iteration, the minimum on the one-dimensional subspace is estimated using Least Square Regression with the posterior mean of the GP. This implies that as more measurements are made on the quantum computer and incorporated into the surrogate model, the GP will provide increasingly accurate estimates.
@@
These modifications distinguish the proposed algorithm from NFT, where the minimum estimation on the subspace depends solely on the previous best point and the fixed shifts for the new measurements. Both modifications aim to make EMICoRe a more flexible and resilient algorithm in the presence of quantum noise. The EMICoRe algorithm starts with $T_{\textrm{NFT}}$ steps of the baseline NFT algorithm to gather sufficient measurements for instantiating an informative GP that can make predictions. After this initial phase, the EMICoRe algorithm begins learning the shifts at each step, using the previous measurements and storing the new measurements for future predictions. For a detailed description of the EMICoRe algorithm and its subroutine, we refer to the original paper [9].
@@
Figure 2 visualizes the algorithmic procedure of the full EMICoRe algorithm. The objective function $E^{\ast}(\boldsymbol{\theta})$ (true energy) is represented by the red lines, while the red shaded area indicates the standard deviation of a measurement on the quantum computer. The blue lines indicate the predictive mean of the GP $\mu_{\boldsymbol{\Theta}}(\cdot)$, while the shaded area denotes its predictive uncertainty $\sqrt{s_{\boldsymbol{\Theta}}(\cdot,\cdot)}$, i.e., one standard deviation. The figure should be read from top left to bottom right following the black arrow.
@@
The top left plot shows the optimization in the direction $\theta_{d}$ for $d\in\{1,\ldots,D\}$ [^2] Note the direction to be optimized is chosen sequentially until $d=D$, after which the algorithm restarts from $d=0$.. The quantity ${\boldsymbol{\hat{\theta}}}^{(t)}$ denotes the set of parameters which give the lowest energy $E({\boldsymbol{\hat{\theta}}}^{(t)})$ from the GP after the $t$-th optimization step. Note that at this point, the GP has access to a dataset of previous measurements $\{\boldsymbol{\Theta}^{(t)},\boldsymbol{\tilde{E}}^{(t)}\}$. This explains why the blue line, along with the corresponding uncertainty, provides a good prediction of the red curve along direction $d$, even before any optimization has been performed.
@@
The EMICoRe subroutine is now responsible for finding the promising points to be measured next, i.e., $\boldsymbol{\Theta}^{\prime}$ in Eq. 7 with the shifts $\alpha$ being different for $\theta_{1}^{\prime}$ and $\theta_{2}^{\prime}$. First, the subroutine divides the subspace into a grid of equally spaced points, as represented by the grid lines at the bottom of the second plot. Second, the subroutine prepares all possible combinations of candidate pairs $\boldsymbol{\Theta}^{\prime}_{k}=\{\boldsymbol{\theta}^{\prime}_{k,2},\boldsymbol{\theta}^{\prime}_{k,2}\}$. These points are added to the base-GP in pairs, allowing a discretized CoRe to be computed for each pair (orange vertical lines). The subroutine then computes the acquisition function for each candidate pair $\boldsymbol{\Theta}^{\prime}_{k}$ using the respective CoRe. The pair that maximizes the acquisition function, as described in Eq. 14, is returned as $\boldsymbol{\Theta}^{\prime}$. According to the algorithm, this pair represents the points that are the most promising to be observed next. The fifth plot displays the measurements on the quantum computer at the suggested points, indicated by the blue crosses. The GP is finally updated using the new measurements—note the change of the blue curve and the shaded area in the plot. Finally, the GP’s posterior mean is computed for three equidistant points on the one-dimensional subspace. These points are then used to fit a cosine function, according to the functional form of the subspace given by Eq. 6. This fitted cosine function is then used to obtain the analytic minimum of the one-dimensional subspace.
@@
We recall that, due to the special form of the VQE-kernel and the design of the EMICoRe algorithm, the function sampled from the GP—i.e, the posterior mean—follows the same functional form as in Eq. 4. Therefore, fitting a sinusoidal function to the GP’s posterior is a valid assumption.
@@ 4 Numerical Experiments
In the following, we perform numerical experiments to demonstrate the performance of EMICoRe in the presence of simulated hardware noise. Our experiments focus on the quantum Heisenberg Hamiltonian, a typical Hamiltonian for benchmarking the performance of VQEs,
@@
where the index of the Pauli matrices $\{X_{j},Y_{j},Z_{j}\}$ refers to the qubit that the Pauli gate is applied to, and open boundary conditions are considered. In this work, we choose the couplings to be at the critical point with $(J_{X},J_{Y},J_{Z})=(-1,0,0)$ and $(h_{X},h_{Y},h_{Z})=(0,0,-1)$. The resulting Hamiltonian is also known as the Ising Hamiltonian at criticality.
In order to assess the performance of the VQE during optimization, we mainly consider two metrics: energy and fidelity. The former is the lowest energy reached by the VQE during the optimization process, i.e., the expectation value of the Hamiltonian with respect to the variational state. The latter measures the overlap between the variational state and the true ground state obtained with exact diagonalization. We note that the true ground-state energy $\underline{E}^{\ast}=\langle\psi_{\mathrm{GS}}|H|\psi_{\mathrm{GS}}\rangle$, along with with the corresponding true ground-state wave function $|\psi_{\mathrm{GS}}\rangle$, can be obtained in our case, as the size of the systems we address is sufficiently small. This is important as it provides access to a ground truth for benchmarking different algorithms, such as NFT versus EMICoRe. We refer to [9] for more details on the metrics we used. To test the stability of the optimization with respect to initialization, we initialize several independent runs with different seeds. Specifically, we use $50$ runs for the experiments without hardware noise and $10$ runs for the experiments with hardware noise, see Section 4.1 for more details. In all plots, we report the median (50th percentile) as a solid line, with the 25th and 75th percentiles as the lower and upper bounds, respectively, represented by shaded regions around the solid line. The results are plotted as a function of the total number of measurements performed on the quantum computer.
For all simulations, we use an $L$-layered and fully connected Efficient SU(2) circuit with $D=2(L+1)\cdot Q$ total angular parameters. The classical simulations of the quantum hardware are performed with the Qiskit [16] library, while the Python implementation of the EMICoRe algorithm has been adapted from the original code available on GitHub [17].
All the parameters for EMICoRe are identical to those chosen for the simulation in Fig. 7 of the original paper [9]. As noted in Footnote 1, we used $\alpha=\pi/3$ as the shift for NFT. Moreover, in contrast to the original NFT paper [7], we did not include a stabilization step, as our results indicated that its use often led to poorer performance. Further investigation of this phenomenon is left to future work.
@@
Figure 3: Energy (left) and fidelity (right) for NFT (green) and EMICoRe (red). We show results for the critical Ising Hamiltonian using $Q=5$ qubits and $L=3$ layers, for an Efficient SU(2) quantum circuit, and $N_{\textrm{shots}}=1024$ measurement shots. The top row displays runs with simulated hardware noise and no error mitigation, while the second and third rows present results with TREX and ZNE error mitigation schemes, respectively. The median (solid line) and the $25^{\mathrm{th}}$ and $75^{\mathrm{th}}$ percentiles (colored regions surrounding the median) are obtained with $50$ independent seeded runs for experiments in Fig. 3(a) and with $10$ runs for Figs. 3(b) and 3(c). A density of the trial distribution at the end of the optimization is shown right next to each plot.